# Spacetime games subsume causal contextuality scenarios

### v3

@@ Abstract
We show that causal contextuality scenarios are equivalent to a subset of the formerly published spacetime games, which generalize game theory to decisions arbitrarily located in Minkowski spacetime. This insight leads to certain constructs and proofs being shorter, simpler, and more intuitive when expressed in the spacetime game framework than in the causal contextuality scenario framework. This shows that the insights of both frameworks taken together can contribute positively to advancing the field of quantum foundations.
@@ 1 Introduction
  • Einsteinian locality: no element of information propagates faster than light.
  • Measurement independence: the choice of a measurement setting, by the experimenter, is exogeneous and, as such, independent of anything not in its future light cone.
Since there is experimental evidence [3] that Bell inequalities are broken by Nature, it follows that one of these two assumptions does not hold. Most of the quantum physics literature so far has focused on assuming measurement independence and dropping Einsteinian locality. Little, but not nothing, has been published on the second possibility, namely, that measurement independence does not hold.
Perhaps this is related to the fact that this assumption of measurement independence is also commonly called “free choice” or “free will” and can lead to strong or emotional philosophical opinions on the matter: we are scientists, but we are also human beings. Yet, it is so that measurement independence is an assumption that can be formalized mathematically in terms of the statistical independence and/or the counterfactual independence (the two are distinct but related in subtle ways [4]) of random variables, and the scientific approach dictates that we must consider the possibility that measurement independence does not hold with as much reason as the possibility that it does. The risk of failing to do so is that an extension theory exists but remains undiscovered because we do not commit any resources to finding it.
One of the most interesting developments in the quantum foundations is the realization that many surprising and disturbing aspects of quantum theory can be abstracted away from physical experiments, and defined purely in terms of which experiments a physicist decides to carry out, and which outcomes they observe. Furthermore, one does not need infinite lists of choices: EPR setups are already interesting with two agents, two settings per agent, and two outcomes per agent and setting (2-2-2). We demonstrated [5] that one can model and view an experimental setup as a game between observers and the universe, Nature being an economic agent that minimizes action (principle of least action). Baczyk and Fourny [6] have given a qualitative argument that the Nash equilibrium [7] paradigm, which has been the mainstream framework of game theory for 75 years, corresponds to the conditions that lead to Bell inequalities:
@@
  • measurement independence corresponds to unilateral deviations;
@@
Subsequent work by Abramsky et al [8] builds on our idea that quantum experimental setups can be seen as games between observers and the universe and models quantum experiments as causal contextuality scenarios. In this paper, we show that our framework subsumes causal contextuality scenarios, i.e., that causal contextuality scenarios are equivalent to a subclass of spacetime games. We provide the formal mapping in both directions. We conclude by showing that the bridge between the two formalisms is mutually beneficial and leads to a clearer understanding of a decision-theoretical view of quantum physics. In particular, we discuss how certain seemingly complex constructs, reasonings, or proofs involving causal contextuality scenarios become one-line arguments in the spacetime game framework.
@@ 2.2 Imperfect information and the dashed lines
Imperfect information is visualized with dashed lines, which group nodes into equivalence classes called information sets. The meaning of these dashed lines is that the agent who makes the decision at these nodes does not know at which one of these nodes they are making their decision: they are not fully informed about the decisions made at ancestor nodes [^2] These ancestor nodes need not all be in the past, they can also be spacelike-separated in the spacetime game. Since information cannot propagate faster than light, it is not possible to be informed about decisions made at spacelike-separated locations. in the extensive form. This is why it is called imperfect information. Concretely, in Figure 2, Bob is not informed of Alice’s decision ($\{X\}$ or $\{Y\}$), and this is why his two nodes are connected with a dashed line. This has to do with relativity theory: although Alice may have already made her decision, the information has not reached Bob yet because his decision is spacelike-separated from Alice’s, as can be seen in the spacelike form of Figure 1, which is a causal dependency graph in the sense of relativity theory.
@@ 2.3 Imperfect information, locality, and non-contextuality
The semantics of imperfect information and dashed lines have a subtle difference depending on whether one looks at the spacetime game or its extensive form. In a spacetime game with imperfect information, dashed lines physically represent non-contextuality, connecting identical decisions made in different contexts, whereas in a game in extensive form with imperfect information (at least those built from a spacetime game), dashed lines represent either locality or non-contextuality. The spacetime game framework makes explicit that locality is a special case of non-contextuality: in local scenarios, the spacetime game has perfect information and the extensive form has imperfect information, while in more general non-contextuality scenarios, both of them have imperfect information.
Figure 2: The game in extensive form with imperfect information that corresponds to the spacetime shown in Figure 1. Alice and Bob have each one information set (A and B) for their choice of measurement setting. Alfred has four information sets X, Y, W, and Z, one for each measurement. For scenarios of interest to the quantum foundations, i.e., exhibiting locality or non-contextuality, the game in extensive form will typically have imperfect information, regardless of whether the spacetime game has perfect information (for locality) or imperfect information (for more general cases of non-contextuality).
@@ 2.4 Causal contextuality scenarios
Subsequent work by Abramsky et al [8] suggests modeling a subclass of such games, which they call causal contextuality scenarios, as strategy presheaves. As we show in this paper, when natural covers are taken, causal contextuality scenarios are isomorphic to a subclass of spacetime games, which we call alternating spacetime games. This means that spacetime games recover causal contextuality scenarios as a special case, and by transitivity (based on the claims made in [8]), that they also recover the framework of Gogioso and Pinzani as a special case [13], as well as Hardy models [14], PR boxes [15], and GHZ models [16].
We also show that a flattened definition of the strategy presheaf based on reduced strategic forms [12] emerges naturally and that the sheaf property holds, and that temporal contextuality scenarios are subsumed by games in extensive form with perfect information.
3 Spacetime games
3.1 Definition
To define the mapping, we now give a formal definition of spacetime games. It is equivalent to that previously published [5] [10], but in a more compact form (see also [6]). In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn [12] and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information. [^3] This is because merging nodes into information sets in the spacetime game translates into merging the corresponding information sets in the extensive form. Spacetime games with perfect information allow for the study of locality, while spacetime games with imperfect information allow for the more general study of non-contextuality.
@@ 3.1 Definition
  • $\mathcal{Z}\in(\mathcal{A}\cup\{\bot\})^{\mathcal{I}}$ is a set of outcomes, also called complete histories, which are partial functions from information sets to actions. They must obey some constraints implied by the game structure. These constraints are defined further down in Section 3.4. The domain of definition, also called path, of an outcome is the set of information sets not mapped to $\bot$, which is the symbol used for unassigned actions.
@@
The payoffs associated with the outcomes are part of a game’s semantics but are not the focus of the present paper, which focuses on the game structure. They are also not part of any causal contextuality scenario and are ignored in the bijective mapping.
3.2 Further examples
@@
Since locality is a special case of non-contextuality, the Bell experiment can also be represented with a single player picking a context with two measurements, in a cyclic manner (XW, WY, YZ, ZX). The corresponding game is shown in Figure 7. The corresponding game in extensive form is shown in Figure 8 and it is almost the same as the game shown in Figure 2, with the observer players merged into one, which amounts to looking at the Cartesian product of the Hilbert spaces.
@@
3.3 Notations
In the remainder of this paper, whenever we refer to a game $\mathcal{G}$, we will also implicitly use the above notations ($\mathcal{N},\mathcal{R}$, $\mathcal{P}$, $\mathcal{A}$, $\rho$, $\chi$, $\sigma$, $\mathcal{I}$) without explicitly defining them every time. This is because we only ever consider one game at a time, and it eases the read.
@@ 3.3 Notations
We also use a slight abuse of notation whenever all nodes in the same information set share a common property: for example, $\rho(i)$ and $\chi(i)$ where $i$ is an information set. We also use, for convenience, a function $\iota\in\mathcal{I}^{\mathcal{N}}$ that maps each node to the information set it belongs to.
@@ 3.4 Histories
In [10], we use the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” [17] [^4] The idea of a decision happening only if a certain causal past is secured, and otherwise not, can be traced back to Dupuy [18][19], and we retain his terminology “causal bridge”. of a node $N$ to specifically designate the predecessors of $N$ and the labels on the edges from the predecessors to $N$. The two differ in that the contingency coordinates involve the entire ancestry, while the causal bridge only involves (directly connected, one level up) predecessors.
@@
$\textsf{domain}(z)=\{i\in\mathcal{I},z(i)\neq\bot\}$
@@ 4 Results
We now formally show that, for a subclass of two-player spacetime games (which we call “alternating”), we can build an equivalent causal contextuality scenario. We also show that, in the other direction, we can convert back causal contextuality scenarios to two-player spacetime games.
Thus, spacetime games and causal contextuality scenarios stand in bijection for a broad class of experimental protocols, which can be either characterized as alternating spacetime games or as causal contextuality scenarios with “good covers” (see open question 7.3 by Abramsky et al [8]), whereby we suggest that a cover is “good” is it corresponds to the natural cover implied by a spacetime game structure.
@@ 5. BOB-A
    $\forall N\in\rho^{-1}(\text{Bob}),\forall a\in\chi(N),\exists M\in\mathcal{N},\sigma(N,M)=a$
@@ 6. BA1
    $\forall N\in\mathcal{N},\rho(N)=\text{Alfred}\implies\exists!P\in\mathcal{N},P\smile N$
@@ 7. BA2
    $\forall N\in\rho^{-1}(\text{Bob}),\forall S,T\in\textsf{successors}(N),\sigma(N,S)=\sigma(N,T)\implies\iota(S)\neq\iota(T)$
@@ 8. AB1
    $\forall i\in\rho^{-1}(\text{Alfred}),\forall N,M\in i,\textsf{successors}(N)=\textsf{successors}(M)$
    $\forall i\in\rho^{-1}(\text{Alfred}),\forall N,M\in i,\forall T\in\textsf{successors}(i),\sigma(M,T)=\sigma(N,T)$

### v4

@@ Abstract
We show that a category of causal contextuality scenarios with no cycles, unique causal bridges, and causally secured covers is equivalent to a category containing a subclass of the formerly published spacetime games, which generalize game theory to decisions arbitrarily located in Minkowski spacetime. This insight leads to certain constructs and proofs being shorter, simpler, and more intuitive when expressed in the spacetime game framework than in the causal contextuality scenario framework. The equivalence of categories and the modular structure of causal contextuality theory also implies that it is possible to build (pure) strategy sheaves, mixed strategy presheaves (equivalent to distribution presheaves) and empirical models on top of spacetime games: the obstruction to a global section in the presence of contextuality corresponds to the non-existence of a mixed strategy in the sense of the Nash game theory framework. This shows that the insights of both frameworks taken together can contribute positively to advancing the field of quantum foundations.
@@ 1 Introduction
  • Freedom of choice: the choice of a measurement setting is independent of anything that could not have been caused by it. This implies [3] (i) measurement independence: the choice of a measurement setting, by the experimenter, is exogeneous and, as such, independent of a hypothetical hidden variable, and (ii) parameter independence: a measurement outcome does not depend on a remote choice of measurement setting.
  • Local determinism [^2] Bell in fact derives this assumption from parameter independence and of the application of the Born rule, which gives a probability of one when both settings are identical.: a measurement outcome is completely determined by what happened in its past light cone.
Since there is experimental evidence [4] that Bell inequalities are broken by Nature, it follows that one of these two assumptions does not hold. Most of the quantum physics literature so far has focused on assuming freedom of choice and dropping local determinism. Little, but not nothing, has been published on the second possibility, namely, that freedom of choice does not hold.
Perhaps this is related to its name: “free choice” or “free will” [5], which can lead to strong or emotional philosophical opinions on the matter: we are scientists, but we are also human beings. Yet, it is so that freedom of choice is an assumption that can be formalized mathematically in terms of the statistical independence and/or the counterfactual independence (the two are distinct but related in subtle ways [6]) of random variables, and the scientific approach dictates that we must consider the possibility that freedom of choice does not hold with as much reason as the possibility that it does. The risk of failing to do so is that an extension theory exists but remains undiscovered because we do not commit any resources to finding it.
One of the most interesting developments in the quantum foundations is the realization that many surprising and disturbing aspects of quantum theory can be abstracted away from physical experiments, and defined purely in terms of which experiments a physicist decides to carry out, and which outcomes they observe. Furthermore, one does not need infinite lists of choices: EPR setups are already interesting with two agents, two settings per agent, and two outcomes per agent and setting (2-2-2). In 2019, we demonstrated [7] that one can model and view an experimental setup as a game between observers and the universe, Nature being an economic agent that minimizes action (principle of least action).
Later, Baczyk and Fourny [8] gave a qualitative argument that the Nash equilibrium [9] paradigm, which has been the mainstream framework of game theory for 75 years, corresponds to the conditions that lead to Bell inequalities:
@@
  • freedom of choice corresponds to unilateral deviations;
@@
Not being aware of our previous work [7][10] of 2019 and 2020 as an antecedent, Abramsky et al. [11] independently rediscovered in 2024 the idea that quantum experimental setups spread in spacetime can be seen as games between observers and the universe. They adapt a previous paper of 2011 [12], which we consider a true masterpiece of mathematics, on sheaf-theoretic structures from the flat case to the causal case. The central, original contribution of [11] is that empirical models can be generalized to the causal case in such a way that non-local or more generally contextual quantum experiments exhibit obstructions to a global section, that is, also when measurements causally dependent on each other. This is done by composing a distribution functor with a strategy presheaf instead of the event sheaf. This provides a considerably stronger and timely quantitative counterpart to our qualitative argument [8].
In this paper, we show that our framework is expressive enough to capture the game structures underlying [11], called causal contextuality scenarios, under some assumptions [^3] In future work, we plan to generalize the equivalence of categories by relaxing some of these assumptions. In this paper we decided to focus on this already generic case.: no cycles, unique causal bridges, and causally secured covers. Formally, under these assumptions, causal contextuality scenarios are equivalent, as a category, to a subcategory of spacetime games that we call alternative spacetime games. We provide the corresponding witness functor.
We conclude by showing that the bridge between the two formalisms is mutually beneficial and leads to a clearer understanding of a decision-theoretical view of quantum physics. In particular, we discuss how certain seemingly complex constructs, reasonings, or proofs involving causal contextuality scenarios become one-line arguments in the spacetime game framework. This includes a short argument that, under the assumptions made in this paper, the strategy presheaf is a sheaf since strategies can be constructed as sets of events using the strategic form [13] similar to the flat case. Likewise, in the opposite direction, the major contribution of [11] on empirical models, thanks to its modular functor-based structure, can be applied on top of alternating spacetime games, since they are structurally equivalent to the causal contextuality scenarios fulfilling our three assumptions.
@@ 2.2 Imperfect information and the dashed lines
Imperfect information is visualized with dashed lines, which group nodes into equivalence classes called information sets. The meaning of these dashed lines is that the agent who makes the decision at these nodes does not know at which one of these nodes they are making their decision: they are not fully informed about the decisions made at ancestor nodes [^4] These ancestor nodes need not all be in the past, they can also be spacelike-separated in the spacetime game. Since information cannot propagate faster than light, it is not possible to be informed about decisions made at spacelike-separated locations. in the extensive form. This is why it is called imperfect information. Concretely, in Figure 2, Bob is not informed of Alice’s decision ($\{X\}$ or $\{Y\}$), and this is why his two nodes are connected with a dashed line. This has to do with relativity theory: although Alice may have already made her decision, the information has not reached Bob yet because his decision is spacelike-separated from Alice’s, as can be seen in the spacelike form of Figure 1, which is a causal dependency graph in the sense of relativity theory. Thus, this corresponds to a form of no-signaling.
@@ 2.3 Imperfect information, locality, and non-contextuality
The semantics of imperfect information and dashed lines have a subtle difference depending on whether one looks at the spacetime game or its extensive form. In a spacetime game with imperfect information, dashed lines physically represent non-contextuality, connecting identical decisions made in different contexts, whereas in a game in extensive form with imperfect information (at least those built from a spacetime game), dashed lines represent either locality or non-contextuality. The spacetime game framework makes explicit that non-locality is a special case of contextuality: in non-local scenarios, the spacetime game has perfect information and the extensive form has imperfect information, while in more general contextuality scenarios, both of them have imperfect information.
Figure 2: The game in extensive form with imperfect information that corresponds to the spacetime shown in Figure 1. Alice and Bob have each one information set (A and B) for their choice of measurement setting. Alfred has four information sets X, Y, W, and Z, one for each measurement. For scenarios of interest to the quantum foundations, i.e., exhibiting non-locality or contextuality, the game in extensive form will typically have imperfect information, regardless of whether the spacetime game has perfect information (for the study of non-locality) or imperfect information (for the more general study of contextuality).
@@ 2.4 Causal contextuality scenarios
Subsequent work by Abramsky et al. [11] suggests modeling a subclass of such games, which they call causal contextuality scenarios, in terms of strategy presheaves. Thanks to a carefully designed modular construction based on a composition of the strategy presheaf with functors, a distribution presheaf is built, with (unsigned or signed) probabilistic or possibilistic semantics, depending on the selection of an underlying semi-ring. Then, contextuality can be witnessed by the obstruction to the existence of a global section in the distribution presheaf, given a set of compatible local distributions with respect to some cover.
The work [11] is itself an adaptation of earlier work on the flat case [12], which brings causality into the picture.
As we show in this paper, certain causal contextuality scenarios—meaning, the underlying structure giving birth to the strategy presheaf—are equivalent to a subclass of spacetime games, which we call alternating spacetime games. This means that spacetime games recover causal contextuality scenarios as a special case. By transitivity (based on the claims made in [11]), this means that by applying the functor connecting the associated strategy presheaf to the distribution presheaf, they can be used to recover the framework of Gogioso and Pinzani as a special case [16], as well as Hardy models [17], PR boxes [18], and GHZ models [19].
We also show that a flattened definition of the strategy presheaf based on strategic forms [13] emerges naturally and that the sheaf property holds555for the strategy presheaf, not for the distribution presheaf, and that temporal contextuality scenarios are subsumed by games in extensive form with perfect information.
2.5 Spacetime games
2.5.1 Definition
We now give a formal definition of spacetime games. It is equivalent to that previously published [7] [10], but in a more compact form (see also [8]). In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn [13] and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information [^6] This is because merging nodes into information sets in the spacetime game translates into merging the corresponding information sets in the extensive form. Spacetime games with perfect information allow for the study of non-locality, while spacetime games with imperfect information allow for the more general study of contextuality.
@@ 2.5.1 Definition
  • $\mathcal{Z}\in(\mathcal{A}\cup\{\bot\})^{\mathcal{I}}$ is a set of outcomes, also called complete histories, which are partial functions from information sets to actions. They must obey some constraints implied by the game structure. These constraints are defined further down in Section 2.5.4. The domain of definition, also called path, of an outcome is the set of information sets not mapped to $\bot$, which is the symbol used for unassigned actions.
@@
The payoffs associated with the outcomes are part of a game’s semantics but are not the focus of the present paper, which focuses on the game structure. They are also not part of any causal contextuality scenario and are ignored for the purpose of the equivalence of categories [^7] In the sense that games with the same structure and different payoffs are isomorphic..
2.5.2 Further examples
@@
Since non-locality is a special case of contextuality, the Bell experiment can also be represented with a single player picking a context with two measurements, in a cyclic manner (XW, WY, YZ, ZX). The corresponding game is shown in Figure 7. The corresponding game in extensive form is shown in Figure 8 and it is almost the same as the game shown in Figure 2, with the observer players merged into one, which amounts to looking at the Cartesian product of the Hilbert spaces.
@@
Figure 9 shows a game that corresponds to the GP example given in [11]. The first observer Alice chooses a measurement X or Y and obtains a result. Then, the second observer chooses a measurement Z or W and obtains a result. On the spacetime form of the game (which in this case is also the extensive form because it is a temporal contextuality scenario), it looks as if the outcome of Bob’s measurement (chosen by Alfred) depends not only on Alice’s choice of measurement but also on the outcome obtained by Alice. However, if one computes the reduced strategic form of the game [13], the inconsistent combinations (e.g., $X=0$ and $Z_{X,1}=0$), which are never activated, get eliminated so that Alfred’s reduced strategy only needs to involve four choices, not eight. This is consistent with [11]’s statement that Alfred’s strategy need only each combination of Alice’s and Bob’s choices to 0 or 1.
Figure 9: A spacetime with perfect information corresponding to a GP scenario.
Figure 10 shows a game [^8] This scenario does not have unique causal bridges and is thus not included in this paper’s categorical equivalence, however an equivalent spacetime game structure still exists as shown here. We plan to later show an extended categorial equivalence also for scenarios with multiple causal bridges. corresponding to an other example given in in [11], which is the implementation of an OR gate on top of a GHZ scenario (Anders-Browne). The choice by Alice of the two inputs were flattened as a cyclic system of rank 4, but could also be expressed with two spacelike-separated agents as shown on Figure 1 (which is equivalent to the cyclic system of rank 4 shown on Figure 7).
Figure 10: A spacetime with perfect information corresponding to the implementation of an OR gate on top of a GHZ scenario. For ease of read, information sets are marked with colors instead of dashed lines. Also, also for the ease of read, the context labels were replaced with dashed lines. The labels can be deduced from the local simplicial complex that arises at each node played by Alice and Bob – another nice visual feature of the spacetime game visualization.
2.5.3 Notations
In the remainder of this paper, whenever we refer to a game $\mathcal{G}$, we will also implicitly use the above notations ($\mathcal{N},\mathcal{R}$, $\mathcal{P}$, $\mathcal{A}$, $\rho$, $\chi$, $\sigma$, $\mathcal{I},u$) without explicitly defining them every time. This is because we only ever consider one game at a time, and it eases the read. When a second game is considered, we prime the letters: a game $\mathcal{G^{\prime}}$, we will also implicitly use ($\mathcal{N^{\prime}},\mathcal{R^{\prime}}$, $\mathcal{P^{\prime}}$, $\mathcal{A^{\prime}}$, $\rho^{\prime}$, $\chi^{\prime}$, $\sigma^{\prime}$, $\mathcal{I}^{\prime}$, u’) without explicitly defining them every time.
For any player $p\in\mathcal{P}$, we write $\mathcal{N}_{p}=\rho^{-1}(p)$. Thus, $\mathcal{N}$ can be organized in a family of disjoint partitions $(\mathcal{N}_{p})_{p\in\mathcal{P}}$.
@@ 2.5.3 Notations
We also use a slight abuse of notation whenever all nodes in the same information set share a common property: for example, $\rho(i)$ and $\chi(i)$ where $i$ is an information set would normally denote a singleton set, but we use them to denote the single element in that singleton set instead, for convenience. Accordingly, for any player $p\in\mathcal{P}$, we write $\mathcal{I}_{p}=\{i\in\mathcal{I}|\rho(i)=p\}$. Thus, $\mathcal{I}$ can be organized in a family of disjoint partitions $(\mathcal{I}_{p})_{p\in\mathcal{I}}$.
We also use, for convenience, a function $\iota\in\mathcal{I}^{\mathcal{N}}$ that maps each node to the information set it belongs to.
@@ 2.5.4 Histories
In [10], we use the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” [20] [^9] We take the terminology from the work of Dupuy [21][22]. of a node $N$ to specifically designate the predecessors of $N$ and the labels on the edges from the predecessors to $N$. The two differ in that the contingency coordinates involve the entire ancestry, while the causal bridge only involves (directly connected, one level up) predecessors.
@@
$\textsf{support}(z)=\{i\in\mathcal{I},z(i)\neq\bot\}$
@@ 3 Results
We will now build a category corresponding to a subclass of two-player spacetime games (which we call “alternating”) as well as a category of causal contextuality scenarios with no cycles, unique causal bridges and causally secured covers, and show that these two categories are equivalent thanks to a witness functor.
Thus, spacetime games and causal contextuality scenarios are equivalent for a broad class of physically meaningful experimental protocols.
@@ 5. BOB-A
    $\forall t\in\mathcal{N}_{B},\forall C\in\chi(t),\exists n\in\mathcal{N}_{A},\sigma(t,n)=C$
@@ 6. BA1
    $\forall n\in\mathcal{N}_{A},\exists!t\in\mathcal{N}_{B},t\smile n$
@@ 7. BA2
    $\forall t\in\mathcal{N}_{B},\forall n,m\in\textsf{successors}(t),\sigma(t,n)=\sigma(t,m)\implies\iota(n)\neq\iota(m)$
@@ 8. AB1
    $\forall x\in\mathcal{I}_{A},\forall n,m\in x,\textsf{successors}(n)=\textsf{successors}(m)$
    $\forall x\in\mathcal{I}_{A},\forall n,m\in x,\forall u\in\textsf{successors}(x),\sigma(n,u)=\sigma(m,u)$